Your company uses the quadratic model y= -11x^2 + 350x to represent how many units (y) of a new product will be sold (x) weeks after its release. How many units can you expect to sell in week 29?
Answer by Boreal(15235)

y(29)=-11*29^2+350(29)=-9251+10150=899.